OpenStudy (anonymous):

x+2y=3 2x-3y=-8 I keep getting -14,-25

OpenStudy (turingtest):

x+2y=3 (1) 2x-3y=-8 (2) solve 1 for x: x=3-2y into (2) 2(3-2y)-3y=-8=6-7y 7y=14 y=2

OpenStudy (turingtest):

you can get x from that, right?